OpenStudy (anonymous):

how do you multiply/divide signed (+/-) fractions? (3/10)(-5/9)

OpenStudy (anonymous):

/

OpenStudy (anonymous):

for divison

OpenStudy (anonymous):

@pink33 that's not even you in your picture

OpenStudy (anonymous):

okay so are you going to help me?

OpenStudy (anonymous):

/

OpenStudy (anonymous):

is for divison

OpenStudy (k12awesomeness):

@sheyennepowell_ Do you mean like negative and positive fractions?

OpenStudy (anonymous):

yea

OpenStudy (k12awesomeness):

Look at the video and it will help you.. They sure helped me when I was learning them:) https://www.khanacademy.org/math/arithmetic/fractions/multiplying_fractions/v/multiplying-negative-and-positive-fractions

OpenStudy (anonymous):

K12? please be aware!!!

OpenStudy (anonymous):

(3/10)(-5/9) = -1*(3/10)(5/9)

OpenStudy (fortytherapper):

Would looking at it like this help? \[\frac{ 3 }{ 10 }\frac{ * }{ * }\frac{ -5 }{ 9 }\]

OpenStudy (anonymous):

If the number of negative signs is odd then the answer result is negative, else positive, ie no negative sign.